OpenStudy (anonymous):

Find the surface area of a cylinder with radius 7 feet and height 16 feet.

OpenStudy (agent0smith):

SA of a cylinder is the area of the 2 bases added together, and the area of the.. outside part of the cylinder, which is circumference times height... Or: \[\Large S.A. = 2*\pi r^2+ 2 \pi r h\] r = 7 feet, h = 16 feet

OpenStudy (anonymous):

ok, so how do we find the circumference with the radius? i understand what you said, but you will need to walk me through it as i missed this day

OpenStudy (anonymous):

\[\bf C=2 \pi r\]

OpenStudy (anonymous):

The cylinder's lateral surface is just a bunch of circumferences piled up on top of each other. Hence you can think of its lateral surface area as the circumference, multiplied by the "height". After that, you must find the area of the circular top/bottom which is evaluted using \(\bf \pi r^2\)

OpenStudy (agent0smith):

Take a can of soup, and wrap a piece of paper around the outside, and cut it so it fits exactly. (or just peel a label off a soup can) When you take the fitted paper off the soup can, notice it'll be a rectangle... with a height equal to the height of the cylinder, and the base of the rectangle, will be the circumference of the circular base.

OpenStudy (anonymous):

@agent0smith nice =D

OpenStudy (agent0smith):

Thanks :) That's how i taught a student i tutored (a girl about 8 or 9) how to understand how to find the lateral SA of a cylinder :) I cut a piece of paper to fit around a can of soup :D